OpenStudy (anonymous):

polar to rectangular solve it in the standard form hyperbola r=8/(2-5 cos (theta))

OpenStudy (anonymous):

\[4y ^{2}-21(x ^{2}-\frac{ 80 }{ 21 }x + \frac{ 40 }{ 21 }-\frac{ 40 }{21})=64\] got stuck at this part

zepdrix (zepdrix):

Hmm before you go to complete the square I wanna see if you missed a negative sign a sec:\[\Large 4y^2-21x^2-80x\quad=\quad64\]\[\Large 4y^2-(21x^2+80x)\quad=\quad64\]

zepdrix (zepdrix):

Did you get that same thing or no? :s

OpenStudy (anonymous):

\[\frac{ (y-0)^{2} }{ (\sqrt{26})^{2} }-\frac{ (x+\frac{ 2\sqrt{210} }{ 21 })^{2} }{ (\sqrt{\frac{ 104 }{ 21 }})^{2} }=1\] final answer

OpenStudy (anonymous):

forgot i factored the negative.....

zepdrix (zepdrix):

:'c

OpenStudy (anonymous):

restart....

OpenStudy (anonymous):

\[4y ^{2}-(-21(x ^{2}+80x))=64\]

OpenStudy (anonymous):

looks right?

zepdrix (zepdrix):

Looks like there is an extra negative.. hmm

OpenStudy (anonymous):

4y2−(21(x2+80x))=64 like this?

zepdrix (zepdrix):

Squaring gave you:\[\Large 4x^2+4y^2\quad=\quad 25x^2+80x+64\]Moving the x's to the left side,\[\Large 4y^2-21x^2-80x\quad=\quad64\]\[\Large 4y^2-21\left(x^2+\frac{80}{21}x\right)\quad=\quad64\]

zepdrix (zepdrix):

When you factor out that 21, don't forget to pull it out of the 80 as well! :) Like you did the first time *

OpenStudy (anonymous):

ok i got that... now completing the square \[4y^2-21(x^2+\frac{ 80 }{ 21 }x +\frac{ 40 }{ 21 }-\frac{ 40 }{ 21 })=64\]

zepdrix (zepdrix):

Woops, you took half of the b value, good... but you forgot to square it.

OpenStudy (anonymous):

\[\frac{ 40 }{ 21 } --> \frac{ 1600 }{ 441 }\]

zepdrix (zepdrix):

I don't think we want to square it. I mean, I think we want to leave the exponent on it. It'll make things a tad easier for us.

OpenStudy (anonymous):

yeah... i wrote the exponent

OpenStudy (anonymous):

\[4y ^{2}-21(x+\frac{ 40 }{ 21 })^{2}=64+21(\frac{ 40 }{ 21 })^2\]

zepdrix (zepdrix):

I think you missed a negative, there are a bunch to keep track of when we move it outside of the brackets.\[\Large 4y^2-21\left(x^2+\frac{80}{21}x+\frac{40^2}{21^2}\color{#CC0033}{-\frac{40^2}{21^2}}\right)\quad=\quad64\]Taking the red term out of the brackets gives us:\[\Large 4y^2-21\left(x^2+\frac{80}{21}x+\frac{40^2}{21^2}\right)\color{#CC0033}{+\frac{40^2}{21}}\quad=\quad64\]Moving to the other side and writing our completed square:\[\Large 4y^2-21\left(x+\frac{40}{21}\right)^2\quad=\quad64\color{#CC0033}{-\frac{40^2}{21}}\]

OpenStudy (anonymous):

then we will have a negative on the right hand side which im again stuck....

zepdrix (zepdrix):

So on the right side, we'll get a common denominator:\[\Large 4y^2-21\left(x+\frac{40}{21}\right)^2\quad=\quad\frac{1344-1600}{21}\] Is that the part you're stuck on? Or earlier than that?

OpenStudy (anonymous):

that is the part

zepdrix (zepdrix):

Subtracting gives us:\[\Large 4y^2-21\left(x+\frac{40}{21}\right)^2\quad=\quad-\frac{256}{21}\]To deal with the negative, we multiply each side by -1,\[\Large 21\left(x+\frac{40}{21}\right)^2-4y^2\quad=\quad\frac{256}{21}\]

OpenStudy (anonymous):

do I multiply everything by \[\frac{21 }{ 256 }\] in order to get 1 on the right side ?

zepdrix (zepdrix):

Ya that sounds right.

OpenStudy (anonymous):

\[\frac{ y^2 }{ (\frac{ 8\sqrt{21} }{ 21 })^{2} }

OpenStudy (anonymous):

... hard to type the final answer

OpenStudy (anonymous):

\[\frac{ y^2 }{ (\frac{ 8\sqrt{21} }{ 21 })^{2} }+ \frac{ x+\frac{ 40^{2} }{ 21 } }{ \frac{ 16 }{ 21 })^{2} }= 1\]

OpenStudy (anonymous):

the ( 40/21) should be (40/21)^2

zepdrix (zepdrix):

Sorry I couldn't get openstudy to work.. \[\Large \frac{y^2}{\left(\cfrac{ \color{red}{16}\sqrt{21}}{21}\right)^{2} }+ \frac{ \color{red}{\left(\color{black}{x+\frac{40}{21}}\right)^2} }{ \left(\cfrac{16}{21}\right)^{2} }= 1\]

zepdrix (zepdrix):

Check out my notes in red, I think you made minor mistake in those spots.

zepdrix (zepdrix):

Oh also, shouldn't the y term be negative? :o

zepdrix (zepdrix):

\[\Large \frac{ \color{red}{\left(\color{black}{x+\frac{40}{21}}\right)^2} }{ \left(\cfrac{16}{21}\right)^{2} }\color{red}{-}\frac{y^2}{\left(\cfrac{ \color{red}{16}\sqrt{21}}{21}\right)^{2}}=1\]

zepdrix (zepdrix):

These problems are painful... Do you have an answer key or some way to check our work?

OpenStudy (anonymous):

nope... its a review question for my coming test on friday

OpenStudy (anonymous):

the answers will be available to us on friday and these will be counted as 5% of our grade for this test

zepdrix (zepdrix):

Oh interesting :O

OpenStudy (anonymous):

why was it \[\frac{ 16\sqrt{21} }{ 21 }\]

zepdrix (zepdrix):

Oh i did something stupid there.. lemme see...

OpenStudy (anonymous):

since it is \[\sqrt{\frac{ 64 }{ 21 }}\]

OpenStudy (anonymous):

rationalized should be\[\frac{ 8\sqrt{21} }{ 21 }\]

zepdrix (zepdrix):

\[\Large \frac{y^2}{\left(\dfrac{16^2}{4\cdot21}\right)}\]The y term is something like that, right?

zepdrix (zepdrix):

\[\Large \frac{4y^2}{\left(\dfrac{16^2}{21}\right)}\quad=\quad \frac{y^2}{\left(\dfrac{16^2}{4\cdot21}\right)}\]

zepdrix (zepdrix):

Oh ok ok ok :) I'm slow lolol Ya I see what you did there :3

OpenStudy (anonymous):

im lost lol...

zepdrix (zepdrix):

\[\Large \frac{y^2}{\left(\dfrac{16^2}{4\cdot21}\right)}\quad=\quad \frac{y^2}{\left(\dfrac{16}{2\sqrt{21}}\right)^2}\]Ok ok now it makes sense :D

zepdrix (zepdrix):

Are you? :3 ah crap!

zepdrix (zepdrix):

\[\Large \frac{ \color{red}{\left(\color{black}{x+\frac{40}{21}}\right)^2} }{ \left(\cfrac{16}{21}\right)^{2} }\color{red}{-}\frac{y^2}{\left(\cfrac{ \color{red}{8}\sqrt{21}}{21}\right)^{2}}=1\] So uhh something like this? +_+

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

gonna have a "great time" graphing this

zepdrix (zepdrix):

lol

OpenStudy (anonymous):

thnx so much by the way

zepdrix (zepdrix):

np c: